OpenStudy (anonymous):

Please help :) Solve: 3(y + 4) - 10 = 66 - 5y

OpenStudy (anonymous):

3y+12-10=66-5y 3y+2=66-5y 2=66-8y -64=-8y 8=y

OpenStudy (anonymous):

thanks, i appreciate it :)